OpenStudy (anonymous):

Does anyone know how to get 3/97 out of this : 1 over z - 1/2z - 1/5z = 10 over 21?

OpenStudy (anonymous):

\[\frac{1}{z}-\frac{1}{2z}-\frac{1}{5z}=\frac{10}{21}\]? i was just writing it. jimmyrep will solve it

OpenStudy (anonymous):

can anyone solve it lol

OpenStudy (anonymous):

multiply thru by 210z 210 - 105 - 42 = 100z 63 = 100z z = 0.63 i cant get 3/97 from this!

OpenStudy (anonymous):

i'm rechecking just in case i made an error

OpenStudy (anonymous):

yep its correct gfost - maybe theres a typo?

OpenStudy (anonymous):

hmm...